assume a single taxpayer is taxed at 10%on the first $9275 of taxable income, 15% on the next $28375 of income, and at 25% for the following$53500 of income.what is the average tax rate for that individual if her taxable income is $42000? show work please
Answer #1

Tax on income = taxable income*tax rate

Tax on first $9,275 = $9,275*10% = ($9,275*10)/100 = $92,750/100 = $927.5

Tax on next $28,375 = $28,375*15% = ($28,375*15)/100 = $425,625/100 = $4,256.25

So far we have taxed total taxable income of $9,275+$28,375 = $37,650. Remaining income to be taxed is total taxable income - income already taxed = $42,000 - $37,650 = $4,350

Tax on remaining $4,350 = $4,350*25% = ($4,350*25)/100 = $108,750/100 = $1,087.5

Average tax rate = (Total tax/taxable income)*100

Average tax rate = [($927.5+$4,256.25+$1,087.5)/$42,000]*100 = ($6,271.25/$42,000)*100 = 0.1493*100 = 14.93%